# Early Detection of Patient Deterioration from Real-Time Wearable Monitoring System

### v1

@@ Abstract
Early detection of patient deterioration is crucial for reducing mortality rates. Heart rate data has shown promise in assessing patient health, and wearable devices offer a cost-effective solution for real-time monitoring. However, extracting meaningful insights from diverse heart rate data and handling missing values in wearable device data remain key challenges. To address these challenges, we propose TARL, an innovative approach that models the structural relationships of representative subsequences, known as shapelets, in heart rate time series. TARL creates a shapelet-transition knowledge graph to model shapelet dynamics in heart rate time series, indicating illness progression and potential future changes. We further introduce a transition-aware knowledge embedding to reinforce relationships among shapelets and quantify the impact of missing values, enabling the formulation of comprehensive heart rate representations. These representations capture explanatory structures and predict future heart rate trends, aiding early illness detection. We collaborate with physicians and nurses to gather ICU patient heart rate data from wearables and diagnostic metrics assessing illness severity for evaluating deterioration. Experiments on real-world ICU data demonstrate that TARL achieves both high reliability and early detection. A case study further showcases TARL’s explainable detection process, highlighting its potential as an AI-driven tool to assist clinicians in recognizing early signs of patient deterioration.
@@ 1 Introduction
Identifying patients at high risk of deterioration is a crucial issue in intensive care units (ICUs). ICUs specialize in treating severely ill patients with more resources and a higher nurse-to-patient ratio than general wards. However, patients in ICUs can deteriorate quickly, and unnoticed changes may delay treatments or transfers to other facilities. Deaths occur during transfers from ICUs Vasilevskis et al. (2009), and each hour of delay can increase mortality odds by 3% Churpek et al. (2016). These challenges highlight the necessity for automated systems that aid clinicians in the early identification of patient deterioration, thereby facilitating prompt interventions and decreasing mortality rates.
Previous research Johansen et al. (2013); Shaffer and Ginsberg (2017) has identified patient’s heart rate data as a key indicator of inflammation, which is associated with various chronic diseases in adult ICU patients Barrett et al. (2014). This implies that consistently monitoring heart rate could be promising for the early identification of patient deterioration. Nonetheless, equipment for ICU monitoring can be expensive and difficult to obtain. Wearable devices provide a cheaper option for ongoing vital sign monitoring, particularly in settings where resources are limited. Grand View Research Grand View Research (2023) valued the wearable technology market at USD 61.30 billion in 2022, emphasizing its growing potential to enhance patient care via real-time monitoring and early intervention.
Meanwhile, ensuring that AI models provide justifiable outcomes is crucial for clinical decision-making. AI models need to deliver justifiable outcomes to help clinicians assess patient conditions and guide trusted treatment decisions Tonekaboni et al. (2019). In this context, heart rate time series data offers a rich source of information for identifying critical health patterns. A promising method for examining heart rate time series is shapelet-based analysis, which has shown effectiveness in multiple fields Zou et al. (2018); Zhu and Hill (2021); Li et al. (2022). A Shapelet Ye and Keogh (2009) refers to a time-series subsequence recognized as a characteristic waveform designed to offer clear explanatory insights into time series data.
@@
To ensure that TARL can effectively assist clinicians in clinical practice, we collaborate with physicians and nurses at the hospital to collect IC patient heart rate data using wearable devices. Also, we collect patients’ laboratory test results and diagnostic metrics to estimate Apache II scores (Acute Physiology and Chronic Health Evaluation II)  Knaus et al. (1985), a widely used indicator of illness severity that serves as ground truth for evaluating patient deterioration in this paper. Our methodology is assessed through real-world ICU patient data, measuring deterioration detection accuracy and earliness to ensure its reliability and impact on assisting clinicians. The interdisciplinary research effort integrates AI researchers, clinicians, and nursing staff, using low-cost wearables and an effective model to strengthen hospital monitoring infrastructure. It has the potential to address the needs of frontline clinicians, supports faster interventions for ICU patients, and offers evidence hospitals can use to improve their triage process.
@@ 2 Related Works
Early Time-Series Classification. Generally, early time-series classification methods can be categorized into three groups: prefix-based, model-based, and shapelet-based. Prefix-based methods determine the minimum prefix length (MPL) required for classification by learning from training instances and applying the learned prefix length to test data.  Xing et al. (2012) first introduce MPL and propose ECTS (early classification on time series). Other works focus on estimating the Minimum Required Length (MRL) for early classification Gupta et al. (2020b, 2021). Various issues are discussed in  Wu et al. (2021). Model-based methods use mathematical models to balance earliness and accuracy Lv et al. (2019); Schäfer and Leser (2019). Shapelet-based methods extract key shapelets to enable reliable label prediction for incomplete time series Yan et al. (2020).
@@ 4.2.2 Transition Confidence
As defined in Eq. 1, missing values in wearable device data are imputed using linear interpolation. Subsequently, we estimate the confidence of a data point $x_{p}$ in the imputed heart rate time series $X$. Let $\Delta t_{p,q}$ be the time difference between $x_{p}$ and the nearest non-imputed data point $x_{q}$ in time series $X$. The confidence $\gamma(x_{p})$ of data point $x_{p}$ is defined as:
$\gamma(x_{p})=\begin{cases}1-\frac{\Delta t_{p,q}}{\varphi}&\text{, if }\Delta t_{p,q}<\varphi\\ 0&\text{, otherwise }\end{cases},$ (4)
where $\varphi>0$ is a fixed value. The smaller the time difference $\Delta t_{p,q}$, the higher the confidence of $x_{p}$. Note that $\gamma(x_{p})$ is set to 1 if $x_{p}$ is a non-imputed data point.
@@ 4.2.3 Model Training
where $\mathbf{v}^{t}_{\tau}\in\mathbb{R}^{d}$ is the embedding of relation $\tau$, and $\mathbf{v}^{\prime t}_{i}$ is the aggregated embedding of shapelet $s_{i}$ obtained from Eq. 3. $(\cdot)^{\text{T}}$ denotes the transpose matrix, and $\text{diag}(\cdot)$ represents the diagonal matrix.
@@ 4.3 Time Series Representation Formulation
Utilizing the embeddings derived from the shapelet-transition KG, each heart rate time series can be represented to reflect the patient’s condition changes throughout the entire sequence $X$.
@@ 5.2 Performance Comparison Results
  • Earliness and Balance. For detecting earliness, while TARL does not achieve the best performance, it consistently ranks among the top three across earliness metrics. Moreover, it achieves the highest Effectiveness-Earliness (EE) score, indicating a strong balance between reliability and early detection. Notably, when TARL attains the highest accuracy, recall, F1-score, and F2-score, its average earliness score reaches 0.79, meaning it detects deterioration about 379 minutes (approximately 6 hours) in advance while observing only 30% of a patient’s heart rate time series. This capability is particularly significant for myocardial infarction (MI), a common ICU condition Barrett et al. (2014), where the golden window for treatment is about 6 hours Steinberg et al. (1994). These results highlight TARL as the most reliable method for detecting deterioration within the golden time of treatment, demonstrating its high practicability for real-world applications.
  • Ablation Study. We observe that removing both transition confidence and the attentive mechanism (TARL w/o T.C. and A.M.) results in the worst detection effectiveness across all metrics and the lowest EE score. Removing transition confidence (TARL w/o T.C.) leads to a more significant drop in average earliness score, likely because the impact of missing values is ignored during embedding, weakening reliable shapelet relationships and delaying deterioration detection. On the other hand, removing the attentive mechanism (TARL w/o A.M.) results in similar performance to TARL, but causes a 6% drop in accuracy. This may be because structurally similar shapelets often exhibit similar future changes, and ignoring these relationships may fail to capture meaningful transitions in heart rates. These findings confirm the necessity of both transition confidence and the attentive mechanism in transition-aware embedding learning.
@@ 5.3 Performance on Different Data Missing Rates
To evaluate TARL under different missing data rates, considering the inherent limitations of wearable devices, we vary the missing rate of heart rate data to {20%, 25%, 30%}, where 20% corresponds to the original dataset. The results are shown in Figure 3. We observed only slight performance variations when applying TARL to datasets with different missing rates. A more detailed analysis is provided below.
  • Although TARL’s performance declines slightly as missing rates increase, its accuracy and recall consistently outperform baselines. While Time2Graph achieves the highest precision, its precision drops significantly with more missing data, whereas TARL maintains stable precision, demonstrating its robustness to missing values in wearable data.
@@ 5.4 Parameter Sensitivity Analysis
To analyze the sensitivity of parameters applied in transition-aware embedding learning, we conduct two sets of experiments: one on the weight ($\omega$) balancing neighborhood proximity and relation proximity in Eq. 3, and the other on the embedding dimension and the number of selected key shapelets. Results are shown in Figure 4 and Figure 5, respectively.
  • In Figure 4, we observe that setting $\omega$ to 0.3 allows TARL to achieve the highest F-score and better earliness performance. While F-scores remain relatively stable across different $\omega$ values, earliness performance becomes more sensitive. As $\omega$ increases, the earliness score gap widens, possibly due to ignoring relation proximity during embedding learning, weakening shapelet transition time interval reinforcement. TARL struggles to accurately capture when a patient’s condition is likely to change, effecting the earliness of the detection.
@@ 6 Conclusion
In this paper, we introduce a novel representation framework, TARL, designed for real-time monitoring using wearable devices to enable early detection of illness deterioration by learning patient heart rate representations. We create a shapelet-transition knowledge graph to model heart rate transitions and devise a transition-aware embedding learning method to enhance shapelet relationships and assess missing data impact. With help from physicians and nurses, we gather ICU patient heart rate data from wearables and diagnostic metrics as deterioration indicators. Experimental results show that TARL achieves the best balance between detection effectiveness and earliness, maintaining strong performance even with varying levels of missing data. Significantly, TARL attains an average earliness score of 0.79, which shows its ability to recognize worsening health conditions 6 hours in advance (the golden treatment window for many treatments), highlighting its practicability in actual hospital settings.
@@ Acknowledgments
We extend our sincere gratitude to Prof. Nai-Ying Ko, Dr. Yi-Ting Chung, and Dr. Chang-Chun Chen from the Department of Nursing, and Prof. Yu-Chen Shu from the Department of Mathematics at National Cheng Kung University, for their invaluable support in wearable device data collection, coordination with patients’ families, and access to physiological data from National Cheng Kung University Hospital.

### v2

@@ Abstract
Early detection of patient deterioration is crucial for reducing mortality rates. Heart rate data has shown promise in assessing patient health, and wearable devices offer a cost-effective solution for real-time monitoring. However, extracting meaningful insights from diverse heart rate data and handling missing values in wearable device data remain key challenges. To address these challenges, we propose TARL, an innovative approach that models the structural relationships of representative subsequences, known as shapelets, in heart rate time series. TARL creates a shapelet-transition knowledge graph to model shapelet dynamics in heart rate time series, indicating illness progression and potential future changes. We further introduce a transition-aware knowledge embedding to reinforce relationships among shapelets and quantify the impact of missing values, enabling the formulation of comprehensive heart rate representations. These representations capture explanatory structures and predict future heart rate trends, aiding early illness detection. We collaborate with physicians and nurses to gather ICU patient heart rate data from wearables and diagnostic metrics to assess illness severity and evaluate deterioration. Experiments on real-world ICU data demonstrate that TARL achieves both high reliability and early detection. A case study further showcases TARL’s explainable detection process, highlighting its potential as an AI-driven tool to assist clinicians in recognizing early signs of patient deterioration.
@@ 1 Introduction
Identifying patients at high risk of deterioration is a crucial issue in intensive care units (ICUs). ICUs specialize in treating severely ill patients with more resources and a higher nurse-to-patient ratio than general wards. However, ICU patients can deteriorate rapidly, and unnoticed changes may delay critical treatments or transfers. Each hour of delay can increase mortality odds by 3% Churpek et al. (2016), and deaths may occur during such transitions. These challenges highlight the necessity for automated systems that aid clinicians in the early identification of patient deterioration, thereby facilitating prompt interventions and decreasing mortality rates.
Previous research Shaffer and Ginsberg (2017) has identified patient’s heart rate data as a key indicator of inflammation, which is associated with various chronic diseases in adult ICU patients. This implies that consistently monitoring heart rate could be promising for the early identification of patient deterioration. Nonetheless, equipment for ICU monitoring can be expensive and difficult to obtain. Wearable devices provide a cheaper option for ongoing vital sign monitoring, particularly in settings where resources are limited. Grand View Research Grand View Research (2023) valued the wearable technology market at USD 61.30 billion in 2022, emphasizing its growing potential to enhance patient care via real-time monitoring and early intervention.
Meanwhile, ensuring that AI models provide justifiable outcomes is crucial for clinical decision-making. AI models need to deliver justifiable outcomes to help clinicians assess patient conditions and guide trusted treatment decisions Tonekaboni et al. (2019). In this context, heart rate time series data offers a rich source of information for identifying critical health patterns. A promising method for examining heart rate time series is shapelet-based analysis, which has shown effectiveness in multiple fields Zhu and Hill (2021); Li et al. (2022). A Shapelet Ye and Keogh (2009) refers to a time-series subsequence recognized as a characteristic waveform designed to offer clear explanatory insights into time series data.
@@
To ensure that TARL can effectively assist clinicians in clinical practice, we collaborate with physicians and nurses at the hospital to collect IC patient heart rate data using wearable devices. Also, we collect patients’ laboratory test results and diagnostic metrics to estimate Apache II scores (Acute Physiology and Chronic Health Evaluation II)  Knaus et al. (1985), a widely used indicator of illness severity that serves as ground truth for evaluating patient deterioration in this paper. Our method is assessed through real-world ICU patient data, measuring deterioration detection accuracy and earliness to ensure its reliability and impact on assisting clinicians. The interdisciplinary research effort integrates AI researchers, clinicians, and nursing staff, using low-cost wearables and an effective model to strengthen hospital monitoring infrastructure. It has the potential to address the needs of frontline clinicians, supports faster interventions for ICU patients, and offers evidence hospitals can use to improve their triage process.
@@ 2 Related Works
Early Time-Series Classification. Early time-series classification methods can be categorized into three groups: prefix-based, model-based, and shapelet-based. Prefix-based methods determine the minimum prefix length (MPL) required for classification by learning from training instances and applying the learned prefix length to test data. Xing et al. first introduce MPL and propose ECTS (early classification on time series) Xing et al. (2012). Other works focus on estimating the Minimum Required Length (MRL) for early classification Gupta et al. (2020b, 2021). Various issues are discussed in  Wu et al. (2021). Model-based methods use mathematical models to balance earliness and accuracy Lv et al. (2019); Schäfer and Leser (2019). Shapelet-based methods extract key shapelets to enable reliable label prediction for incomplete time series Yan et al. (2020).
@@ 4.2.2 Transition Confidence
As defined in Def. 1, missing values in wearable device data are imputed using linear interpolation. Subsequently, we estimate the confidence of a data point $x_{p}$ in the imputed heart rate time series $X$. Let $\Delta t_{p,q}$ be the time difference between $x_{p}$ and the nearest non-imputed data point $x_{q}$ in time series $X$. The confidence $\gamma(x_{p})$ of data point $x_{p}$ is defined as:
$\gamma(x_{p})=\Big{(}1-\frac{\Delta t_{p,q}}{\varphi}\Big{)}\cdot\mathbb{I}(\Delta t_{p,q}<\varphi)$ (4)
where $\varphi>0$ is a fixed value, and $\mathbb{I}(\cdot)$ is an indicator function. The smaller the time difference $\Delta t_{p,q}$, the higher the confidence of $x_{p}$. Note that $\gamma(x_{p})$ is set to 1 if $x_{p}$ is a non-imputed data point.
@@ 4.2.3 Model Training
where $\mathbf{v}^{t}_{\tau}\in\mathbb{R}^{d}$ is the relation $\tau$’s embedding, and $\mathbf{v}^{\prime t}_{i}$ is the aggregated embedding of shapelet $s_{i}$ from Eq. 3. $(\cdot)^{\text{T}}$ denotes transpose, and $\text{diag}(\cdot)$ the diagonal matrix.
@@ 4.3 Time Series Representation Formulation
Utilizing the embeddings derived from the shapelet-transition KG, each heart rate time series $X$ can then be represented to reflect the patient’s condition changes.
@@ 5.2 Performance Comparison Results
  • Earliness and Balance. While TARL isn’t the top performer in earliness, it consistently ranks in the top three and achieves the highest Effectiveness-Earliness (EE) score, demonstrating a strong balance between reliability and early detection. Notably, when TARL attains the highest accuracy, recall, F1-score, and F2-score, its average earliness score reaches 0.79, meaning it detects deterioration about 379 minutes (approximately 6 hours) in advance while observing only 30% of a patient’s heart rate time series. This capability is particularly significant for myocardial infarction (MI), a common ICU condition, where the golden window for treatment is about 6 hours Steinberg et al. (1994). These results highlight TARL as the most reliable for detecting deterioration within the golden time of treatment, demonstrating its high practicability for real-world applications.
  • Ablation Study. Removing both transition confidence and the attentive mechanism (TARL w/o T.C. and A.M.) results in the worst detection effectiveness and the lowest EE score. Excluding only transition confidence (TARL w/o T.C.) causes a larger drop in earliness score, likely because the impact of missing values is ignored during embedding, weakening reliable shapelet relationships and delaying deterioration detection. On the other hand, removing the attentive mechanism (TARL w/o A.M.) causes a 6% drop in accuracy. This may be because structurally similar shapelets often exhibit similar future changes, and ignoring these relationships may miss meaningful heart rate transitions. These findings confirm the importance of both transition confidence and attentive mechanism in TARL.
@@ 5.3 Performance on Different Data Missing Rates
To simulate missing data in wearable devices, we vary the missing rate of heart rate data to {20%, 25%, 30%} to evaluate TARL’s performance. As shown in Figure 3, TARL’s performance varies only slightly across rates. A more detailed analysis is provided below.
  • Although TARL’s performance slightly declines with higher missing rates, its accuracy and recall consistently surpass baselines. While Time2Graph achieves the highest precision, its precision drops significantly with more missing data, whereas TARL maintains stable precision, demonstrating its robustness to missing values in wearable data.
@@ 5.4 Parameter Sensitivity Analysis
To analyze parameter sensitivity in TARL, we conduct two experiments: one on the weight ($\omega$) balancing neighborhood and relation proximities in Eq. 3, and the other on the embedding dimension and the number of selected key shapelets. Results are shown in Figure 4 and Figure 5, respectively.
  • As shown in Figure 4, setting $\omega$ to 0.3 yields the best F-score and improved earliness. While F-scores remain relatively stable across $\omega$, earliness performance is more sensitive. As $\omega$ increases, the earliness score gap widens, possibly due to ignoring relation proximity during embedding learning, weakening shapelet transition time interval reinforcement. TARL struggles to accurately capture when a patient’s condition is likely to change, effecting the earliness of the detection.
@@ 6 Conclusion
In this paper, we introduce a novel representation framework, TARL, designed for real-time monitoring using wearable devices to enable early detection of illness deterioration by learning patient heart rate representations. We create a shapelet-transition knowledge graph to model heart rate transitions and devise a transition-aware embedding learning method to enhance shapelet relationships and assess missing data impact. With help from physicians and nurses, we gather ICU patient heart rate data from wearables and diagnostic metrics as deterioration indicators. Experimental results show that TARL achieves the best balance between detection effectiveness and earliness, maintaining strong performance even with varying levels of missing data. A case study also illustrates TARL’s explainable detection process, which has the potential to help clinicians identify early signs of patient deterioration.
@@ Acknowledgments
We sincerely thank Prof. Nai-Ying Ko, Dr. Yi-Ting Chung, and Dr. Chang-Chun Chen from the Department of Nursing, and Prof. Yu-Chen Shu from the Department of Mathematics at National Cheng Kung University (NCKU), for their invaluable support in wearable data collection, patient coordination, and access to physiological data from NCKU Hospital.
In addition, this paper was supported in part by National Science and Technology Council (NSTC), R.O.C., under Contract 113-2221-E-006-203-MY2, 114-2622-8-006-002-TD1 and 113-2634-F-006-001-MBK.